The lifetime of a certain type of automatic machine is normally distributed with mean and standard deviation equals to 3.1 and 1.1 years repspectivelly, use the standard normal table to find P left parenthesis x less than 3.3 right parenthesis
1
Expert's answer
2020-12-11T12:49:37-0500

"X\sim N(3.1, 1.1^2)"

"P(X<3.3)=P(Z<\dfrac{3.3-\mu}{\sigma})"

"=P(Z<\dfrac{3.3-3.1}{1.1})\approx P(Z<0.181818)"

"\approx0.572137"